OpenStudy (samigupta8):

In case of organ pipes Can anyone pls...help me describe the equation for pressure variation for standing waves formed?

OpenStudy (samigupta8):

@priyar, @mayankdevnani,@parthkohli pls help

OpenStudy (samigupta8):

@parthkohli

OpenStudy (mayankdevnani):

you want to learn from me (time wasting) or i'll give you a better site,learn from there

OpenStudy (mayankdevnani):

it is totally physics based site,created by EX FIITJEE PHYSICS HOD,EX RESONANCE PHYSICS HOD,EX NARAYANA HOD,PRESENT ALLEN HOD JAIPUR.

OpenStudy (mayankdevnani):

all are same person....haha

OpenStudy (samigupta8):

ashish arora

OpenStudy (samigupta8):

Pg ki Baat kr re ho

OpenStudy (mayankdevnani):

haan

OpenStudy (samigupta8):

Us me h yeh concept

OpenStudy (mayankdevnani):

lol..tumne saara suspense kharab kr diya

OpenStudy (mayankdevnani):

haan usme h yeh concept

OpenStudy (samigupta8):

Okk...

OpenStudy (samigupta8):

Tysm nyways....

OpenStudy (mayankdevnani):

vaise tumhe yeh pata lagana h ki yeh kaha se aaya :- \[\large \bf \triangle P_{net}=(2 \triangle P_0\cos kx)(\sin \omega t)\]

OpenStudy (mayankdevnani):

where, \[\large \bf A=Amplitude=2 \triangle P_0 \cos kx\]

OpenStudy (mayankdevnani):

ab aa gaya samaj main ?

OpenStudy (samigupta8):

Actually I m confused in dis thing that Pressure form differs displacement wave by π/2 phase so pressure form shud be Asin(wt-kx+π/2)

OpenStudy (samigupta8):

BT in a book it is given that it is Asin(wt-kx) how????

OpenStudy (samigupta8):

Haan exactly wohi Jo likha tha tumne abhi

OpenStudy (mayankdevnani):

\[\large \bf \triangle P=\triangle P_0\sin(\omega t-kx+\phi)\]

OpenStudy (samigupta8):

Main del symbol ni bna paari thi

OpenStudy (mayankdevnani):

same thing

OpenStudy (samigupta8):

Kahan se aaya h yeh

OpenStudy (mayankdevnani):

equation kya ?

OpenStudy (samigupta8):

Yup

OpenStudy (mayankdevnani):

just like SHM

OpenStudy (mayankdevnani):

Remember :- Sound Wave can be expressed in form of PRESSURE DIFFERENCE WAVE.

OpenStudy (samigupta8):

Haan fir

OpenStudy (samigupta8):

Shm kse lg raha h yeh tumhe yeh to wave ki equation h

OpenStudy (samigupta8):

Shm to tbb bnti h jbb stationary waves hoti h

OpenStudy (mayankdevnani):

Hence,in form of displacement ,sound wave is written as :- \[\large \bf \triangle S=S_0\sin(\omega t-kx+\phi+\frac{\pi}{2})\] \[\large \bf \triangle S=S_0\cos(\omega t-kx+\phi)\]

OpenStudy (samigupta8):

Abhi I m talking only about the single equation

OpenStudy (samigupta8):

Haan to yeh to cos ki hui NAA abb N isski jbb stationary wave ki equation likhoge to dono cos ki terms aayegi

OpenStudy (samigupta8):

CosC +CosD se

OpenStudy (mayankdevnani):

oye!

OpenStudy (mayankdevnani):

standing wave hamesha pressure variation main lete h

OpenStudy (mayankdevnani):

displacement variation to string main lete h

OpenStudy (samigupta8):

Arey ni usme b lete h

OpenStudy (mayankdevnani):

https://www.youtube.com/watch?v=HvKej75_csQ

OpenStudy (mayankdevnani):

yaar mujhe kuch nahi samaj main aa raha h ki tum kya kehna chahti ho ?

OpenStudy (mayankdevnani):

tum yeh to nahi kehna chahti ki mujhe standing wave displacement variation main chahiye kyunki question yeh pooch raha h

OpenStudy (samigupta8):

Luk what I m asking is why the pressure variation is of form delPo sin(wt-kx)

OpenStudy (mayankdevnani):

dekho main samjata hoon kyun

OpenStudy (mayankdevnani):

string waves main displacement variation lete h kyunki string main displacement hota h right?

OpenStudy (samigupta8):

Arey isme b hota h displacement variation shi bta ri hu main buk me b h n even I m sure of it...

OpenStudy (mayankdevnani):

arrey mujhe pata h suno to

OpenStudy (samigupta8):

Okkkk

OpenStudy (mayankdevnani):

sound wave=longitudinal wave=pressure difference wave=pressure wave

OpenStudy (mayankdevnani):

isliye hum general form lete h:- del sin(wt-kx)

OpenStudy (samigupta8):

BT pressure variation toh π/2 phase difference me hota h na uske displacement k

OpenStudy (mayankdevnani):

ulta bol rahi ho

OpenStudy (samigupta8):

Kya

OpenStudy (mayankdevnani):

displacement variation toh π/2 phase difference me hota h na uske pressure k

OpenStudy (samigupta8):

Arrey phase difference π/2 h bass....un donon k bich

OpenStudy (mayankdevnani):

haan lekin pressure variation ke respect main

OpenStudy (samigupta8):

Haan to abb equation change ho gyi naa

OpenStudy (samigupta8):

Cos me aa gyi naa

OpenStudy (mayankdevnani):

okay

OpenStudy (mayankdevnani):

then simply add up

OpenStudy (samigupta8):

Haan to do no cos me aayenge maa

OpenStudy (mayankdevnani):

haan

OpenStudy (samigupta8):

Haan to ni h naa woh buk me

OpenStudy (mayankdevnani):

phir mujhe kuch nahi pata

OpenStudy (mayankdevnani):

ab batao main kya kar sakta hoon isme? hukm mere aaka

OpenStudy (samigupta8):

Kuch b ni h hukm....

OpenStudy (mayankdevnani):

phir

OpenStudy (mayankdevnani):

abhi main sonne jaunga

OpenStudy (mayankdevnani):

10-15 min main

OpenStudy (samigupta8):

Okkie